# Widespread disruption of resonant chains during protoplanetary disk dispersal

### v1

@@ 5.2.1 Initial conditions
Figure 14 illustrates the effect of late time inspiral by showing the evolution of a system with the same mass configuration as in Figure 13, with the new, more distant, initial conditions. The inner planet reaches the torque reversal relatively early (at A), and the middle planet catches up to it at B. The systems started wider in this case and so they capture into the 2:1 resonance at this point and the torques on the middle planet drive the pair inwards until the middle planet reaches the torque reversal at C. Eventually, the outer planet gets close enough to encounter the 2:1 resonance with the middle planet at D.
@@ 6 Discussion
The Kepler satellite detects planets via their transits, so we must calculate the geometric probability that each of the planets in our simulations are actually observed. The probability of observing the transit of an individual planet is straightforward, scaling as the inverse of the semi-major axis. The relative frequency of transit of members of multiple planet systems depends on the inclination dispersion within the system. To assess these probabilities, we observe each system from $10^{5}$ random orientations on the sky. For each realisation, the mutual inclinations are drawn from a Rayleigh distribution with an inclination dispersion of $3^{\circ}$, the upper limit consistent with the Kepler statistics (Fang & Margot, 2012). The relative lines of nodes are assumed to be randomly distributed. With these results, we determine the relative frequency of observing all three planets in each system, and the frequency of observing the three different permutations of planetary pairs.
@@
Figure 15 shows the resulting probability of observing different period ratios between neighbouring planets. In the upper panel, we show the observed distribution of Kepler period ratios. We show two samples. The first is the standard compilation, from Fabrycky et al. (2014), shown as a dotted histogram. Since this is now ten years old, and other systems have been confirmed since, we also show, as a solid histogram, those planets identified in the NASA Exoplanet Archive (Akeson et al., 2013) as confirmed, and discovered by Kepler. The two first order resonances (q+1)/q, with q=1 and q=2, are shown as dashed lines, illustrating the well-known result that most nearest neighbour pairs in the Kepler sample are not near a commensurability.
@@
Figure 15: The upper panel shows the observed distribution of nearest neighbour period ratios in the sample of Kepler systems, based on the compilation of Fabrycky et al. (2014) (dotted histogram) and the latest compilation of confirmed planets, discovered by Kepler, from the NASA exoplanet archive. The three panels below that show the outputs of our simulations. The second panel shows the final distributions that started with compact initial conditions and mass ratios drawn from the observed Kepler triple systems, and assuming that all stars the default value of the transition radius $a_{R}$. The third panel shows the results of similar initial conditions, but adding together systems drawn equally from $a_{R}=0.8$, $a_{R}=1.2$ and $a_{R}=1.5$. The bottom panel shows the results from the simulations that begain with more distant, and wider spaced, initial conditions. This means that more systems get trapped into the 2:1 resonance instead of the 3:2 resonances, and the later divergence leads to wider systems. These simulations assume the standard $a_{R}$.
@@
The other panels in Figure 15 show the output from our simulations. The second panel from the top, labelled as ‘Compact’, represents the results from our default simulations (the ones shown as red points in Figure 10). These have been assigned weights according to their geometric probability of observation, as discussed above. The third panel from the top represents the evolution of the same set of initial conditions, but with $a_{R}$ scaled up by a factor of 1.2 relative to the default value given in equation (4). This represents the case of a stronger stellar magnetic fields. The resulting distribution is pretty similar, although the resonant peaks are less pronounced – i.e. a larger fraction of systems moved away from commensurabilities.
@@
Finally, the lower panel shows the results of starting with wide initial conditions (and an $a_{R}$ corresponding to that of the second panel, i.e. all systems have the default $a_{R}$). We see that many more of these systems get captured in the 2:1 resonance and we get a distribution that skews to wider separations than expected in the observations.
@@
A more quantitative comparison of the non-resonant fractions is shown in Figure 16, which compares the cumulative distribution of period ratios between theory and observations.
@@
The black histogram represents the observed distribution drawn from the current Kepler sample (the solid histogram in the upper panel Figure 15). The red histogram represents the default model sample from the second panel, and which reproduces the observed distribution quite accurately. The magenta histogram shows the distribution from the third panel. This one shows too many at lower period ratios and too few at higher period ratios. The blue histogram shows the sample from the lowest panel – those drawn from wide initial conditions. In this case the model clearly produces too few close pairs and even falls below the observed curve at larger periods, indicating a substantial pile-up of systems between period ratios of 2.2 and 3. Overall, we find that the default model does an excellent job of providing a distribution of period ratios broadly consistent with those observed. A Kolmogorov-Smirnov test indicates that the two distributions have a 70% chance of being drawn from the same underlying distribution. The observations can also be reproduced with combinations of the different samples. The cyan histogram shows a model in which 70% of the systems come from the simulations with larger $a_{R}$ and compact initial conditions, along with 30% of the systems that began with wider initial conditions.
@@
Another useful comparison set is the simulations of Izidoro et al. (2017) and Izidoro et al. (2021), the most comprehensive effort to date to compare the migration model with observations in a quantitative fashion. Izidoro et al. (2017) describe a model for the migration of low mass planets in both static and turbulent disks, which decay away on Myr timescales, leaving planets trapped in a chain of mean motion resonances. They furthermore run the simulations forward a further 100 Myr after the dispersal of the gas disk, and track the evolution of systems that undergo dynamical instabilities which can break the resonant chains. The encouraging part of the results of Izidoro et al. (2017) is that the resulting instabilities do indeed produce many non-resonant systems. However, the fraction of resonant systems that remain is still far too high ($\sim 50\%$) and the dynamical instabilties lead to planetary collisions which result in systems too massive compared to the observed systems. The simulations of Izidoro et al. (2021) takes these models further by including a prescription for gas-assisted pebble accretion during the migration phase. Variation in the pebble flux and the formation time of planetesimals result in various levels of dynamical instability. The authors find that it is possible to match the neighbour period ratio distribution of the Kepler planets but, once again, only if one treats the fraction of unstable systems as a free parameter. For some parameter choices, most of the simulated systems do go unstable, but these also, once again, produce systems that are too massive.
@@
Indeed, the distribution is actually quite similar to that which emerges from simulations that model the planetary configurations emerging from in situ assembly. The dotted histogram shows the outcomes from the simulations of Hansen & Murray (2013) for the case of a total solid inventory of $20M_{\oplus}$. The magnetospheric rebound model actually produces systems that are even more compact than the in situ models, despite the fact that the latter are partially sculpted by considerations of dynamical stability, while the spacing of the former are regulated by the freeze-out process of the gaseous disk.
The time at which a given planet freezes out of the migration will also determine the final orbital period of the planet. This will also be affected by the value of $a_{R}$ and hence the stellar magnetic field. To test whether our model values are reasonable, we smulated the same 107 Kepler systems with the default value of $a_{R}$, a version with $a_{R}$ scaled up by a factor 1.2 (discussed above already) and also one with a scaling factor of 0.8. Figure 18 shows the resulting distribution of final planetary locations for all planets summed over these three simulations. In this case we do not adjust the counts for the geometric probabilities because the observations to which we compare have already accounted for selection effects and purport to represent the true underlying distribution. The shaded regions are these corresponding observed trends in planet occurrence for Super-Earths, as measured by Petigura et al. (2018). The simulations provide an excellent match to the observed rise from short periods to a peak in the range of $\sim$ 6–8 days, followed by a drop-off towards larger periods. Immediately exterior to the peak, the simulations also match the slope measured by Petigura et al. (2018) but begin to drop below this at orbital periods $>$20 days. This is a reflection of our restriction to triple systems. Systems with additional planets are expected to further fill in the distribution at longer periods.
@@ 7 Conclusions
The results presented here demonstrate that a detailed treatment of the inner boundary of protoplanetary disks Yu et al. (2023) has a strong influence on the final configurations of planetary systems that migrate through this disk. In particular, these models show the existence of an inner transition region of outwardly directed migration, which causes divergent migration of planets in resonant chains at late times. This results in a substantial fraction of initially resonant chains diverging to the point that their period ratios no longer lie close to first order commensurabilities (although some may still be technically resonant because of their high precession rates). The resulting distribution of planetary neighbour period ratios is a good match to the observed distribution of planets detected by Kepler, and also provides a good match to the observed relative occurrence frequency of observed Super-Earth planets. These results provide an explanation for the long-standing problem that the observed distribution of period ratios showed far fewer resonant systems than expected in the models.
@@
This research has made use of the NASA Exoplanet Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. This research has made use of NASA’s Astrophysics Data System Bibliographic Services.
@@ Appendix A A: Dynamical Equations
The planetary dynamics will be described in terms of the orbital frequencies, $n_{1}$, $n_{2}$ and $n_{3}$, normalised to a fidicual value $n_{0}$, so that $x_{i}=n_{i}/n_{0}$. These correspond to semi-major axis $a_{1}<a_{2}<a_{3}$, so that $\alpha_{12}=a_{1}/a_{2}$ and $\alpha_{23}=a_{2}/a_{3}$. Eccentricities are denoted by $e_{i}$ and the arguments of periastron are designed as $\bar{\omega}_{i}$. We will incorporate effects of the first order resonances $(1+q)/q$ for $q=1$ and $q=2$, between nearest neighbours. The resonant angles are $\phi_{1}=2\lambda_{2}-\lambda_{1}$, $\phi_{2}=3\lambda_{2}-2\lambda_{1}$, $\phi_{3}=2\lambda_{3}-\lambda_{2}$ and $\phi_{4}=3\lambda_{3}-2\lambda_{2}$. We normalise the timescale to $t_{0}=M_{*}/(3n_{0}m_{2})$, so that $\dot{x}_{i}=dx_{i}/d\tau=t_{0}dx_{i}/dt$.
@@
$\dot{x}_{1}$ $=$ $-x_{1}^{2}\alpha_{12}\left[A_{12}\sin\phi_{1}-B_{12}\cos\phi_{1}\right]-2x_{1}^{2}\alpha_{12}\left[C_{12}\sin\phi_{2}-D_{12}\cos\phi_{2}\right]+\frac{3}{2}\frac{x_{1}}{\tau_{a,1}}+3e_{1}^{2}\frac{x_{1}}{\tau_{e,1}}$ (7)
$\dot{x}_{2}$ $=$ $2x_{2}^{2}\frac{m_{1}}{m_{2}}\left[A_{12}\sin\phi_{1}-B_{12}\cos\phi_{1}\right]+3x_{2}^{2}\frac{m_{1}}{m_{2}}\left[C_{12}\sin\phi_{2}-D_{12}\cos\phi_{2}\right]+$ (8)
$-x_{2}^{2}\alpha_{23}\frac{m_{3}}{m_{2}}\left[A_{23}\sin\phi_{3}-B_{23}\cos\phi_{3}\right]-2x_{2}^{2}\alpha_{23}\frac{m_{3}}{m_{2}}\left[C_{23}\sin\phi_{4}-D_{23}\cos\phi_{4}\right]+\frac{3}{2}\frac{x_{1}}{\tau_{a,1}}+3e_{2}^{2}\frac{x_{2}}{\tau_{e,2}}$
$\dot{x}_{3}$ $=$ $2x_{3}^{2}\left[A_{23}\sin\phi_{3}-B_{23}\cos\phi_{3}\right]+3x_{3}^{2}\left[C_{23}\sin\phi_{4}-D_{23}\cos\phi_{4}\right]+\frac{3}{2}\frac{x_{3}}{\tau_{a,3}}+3e_{3}^{2}\frac{x_{3}}{\tau_{e,3}}$ (9)
@@
The eccentricity damping in these equations can often lead to low eccentricities which, in turn, leads to high precession rates. As a result, some of the resonant angles can continue to librate even when the period ratios deviate from the exact commensurability. Our equations represent the expansions in terms of the resonant arguments for the instantaneous osculating value of $\alpha$, and so we describe the functions $f_{j}$ as functions of $\alpha$ and not simply a number at exact commensurability.
@@
$\dot{h}_{1}$ $=$ $-\frac{h_{1}}{\tau_{e,1}}+\frac{\alpha_{12}}{3}f_{1}(\alpha_{12})x_{1}\cos\phi_{1}+\frac{\alpha_{12}}{3}f_{3}(\alpha_{12})x_{1}\cos\phi_{2}$ (21)
$\dot{k}_{1}$ $=$ $-\frac{k_{1}}{\tau_{e,1}}-\frac{\alpha_{12}}{3}f_{1}(\alpha_{12})x_{1}\sin\phi_{1}+\frac{\alpha_{12}}{3}f_{3}(\alpha_{12})x_{1}\sin\phi_{2}$ (22)
$\dot{h}_{2}$ $=$ $-\frac{h_{2}}{\tau_{e,2}}+\frac{1}{3}\frac{m_{1}}{m_{2}}f^{\prime\prime}_{2}(\alpha_{12})x_{2}\cos\phi_{1}+\frac{1}{3}\frac{m_{1}}{m_{2}}f_{4}(\alpha_{12})x_{2}\cos\phi_{2}$ (23)
$+\frac{\alpha_{23}}{3}\frac{m_{3}}{m_{2}}f_{1}(\alpha_{23})\cos\phi_{3}+\frac{\alpha_{23}}{3}\frac{m_{3}}{m_{2}}f_{3}(\alpha_{23})x_{2}\cos\phi_{4}$
$\dot{k}_{2}$ $=$ $-\frac{k_{2}}{\tau_{e,2}}-\frac{1}{3}\frac{m_{1}}{m_{2}}f^{\prime\prime}_{2}(\alpha_{12})x_{2}\sin\phi_{1}-\frac{1}{3}\frac{m_{1}}{m_{2}}f_{4}(\alpha_{12})x_{2}\sin\phi_{2}$ (24)
$-\frac{\alpha_{23}}{3}\frac{m_{3}}{m_{2}}f_{1}(\alpha_{23})x_{2}\sin\phi_{3}+\frac{\alpha_{23}}{3}\frac{m_{3}}{m_{2}}f_{3}(\alpha_{23})x_{2}\sin\phi_{4}$
$\dot{h}_{3}$ $=$ $-\frac{h_{3}}{\tau_{e,3}}+\frac{1}{3}f^{\prime\prime}_{2}(\alpha_{23})x_{3}\cos\phi_{3}+\frac{1}{3}f_{4}(\alpha_{23})x_{2}\cos\phi_{4}$ (25)
$\dot{k}_{3}$ $=$ $-\frac{k_{3}}{\tau_{e,3}}-\frac{1}{3}f^{\prime\prime}_{2}(\alpha_{23})x_{3}\sin\phi_{3}-\frac{1}{3}f_{4}(\alpha_{23})x_{3}\sin\phi_{4}$ (26)
Finally, the above equations depend on the commensurabilities $\phi_{1}=2\lambda_{2}-\lambda_{1}$, $\phi_{2}=3\lambda_{2}-2\lambda_{1}$, $\phi_{3}=2\lambda_{3}-\lambda_{2}$ and $\phi_{4}=3\lambda_{3}-2\lambda_{2}$. The evolution of these variables is given by
$\dot{\phi}_{1}$ $=$ $\frac{1}{\Delta}\left(2x_{2}-x_{1}\right)$ (27)
$\dot{\phi}_{2}$ $=$ $\frac{1}{\Delta}\left(3x_{2}-2x_{1}\right)$ (28)
$\dot{\phi}_{3}$ $=$ $\frac{1}{\Delta}\left(2x_{3}-x_{2}\right)$ (29)
$\dot{\phi}_{4}$ $=$ $\frac{1}{\Delta}\left(3x_{3}-2x_{2}\right)$ (30)
where $\Delta=3m_{2}/M_{*}$.

### v2

@@ 5.2.1 Initial conditions
Figure 13 illustrates the effect of late time inspiral by showing the evolution of a system with the same mass configuration as in the right hand side of Figure 12, with the new, more distant, initial conditions. The inner planet reaches the torque reversal relatively early (at A), and the middle planet catches up to it at B. The systems started wider in this case and so both pairs capture into the 2:1 resonance at this point and the torques on the middle planet drive the pair inwards until the middle planet reaches the torque reversal at C. Eventually, the outer planet gets close enough to encounter the 2:1 resonance with the middle planet at D.
@@ 6 Discussion
The Kepler satellite detects planets via their transits, so we must calculate the geometric probability that each of the planets in our simulations are actually observed. This scales as the inverse of the semi-major axis. The relative frequency of transit of members of multiple planet systems depends on the inclination dispersion within the system. To assess these probabilities, we observe each system from $10^{5}$ random orientations on the sky. For each realisation, the mutual inclinations are drawn from a Rayleigh distribution with an inclination dispersion of $3^{\circ}$, the upper limit consistent with the Kepler statistics (Fang & Margot, 2012). The relative lines of nodes are assumed to be randomly distributed. With these results, we determine the relative frequency of observing all three planets in each system, and the frequency of observing the three different permutations of planetary pairs.
@@
Figure 14 shows the resulting probability of observing different period ratios between neighbouring planets. In the upper panel, we show the observed distribution of Kepler period ratios, using those planets identified in the NASA Exoplanet Archive (Akeson et al., 2013) as confirmed, and discovered by Kepler. The resulting distribution resembles that of the standard compilation, from Fabrycky et al. (2014), but with improved statistics. The two first order resonances (q+1)/q, with q=1 and q=2, are shown as dashed lines, illustrating the well-known result that most nearest neighbour pairs in the Kepler sample are not near a commensurability.
@@
Figure 14: The upper panel shows the observed distribution of nearest neighbour period ratios in the sample of Kepler systems, based on the data from the NASA exoplanet archive, with Poisson error bars. The three panels below that show the outputs of our simulations. The second panel (red histogram) shows the final distributions of the simulations described in § 5.2, which assumes that all stars the default value of the transition radius $a_{R}$. The third panel (magenta histogram) shows the results of similar initial conditions, but using $a_{R}=1.2$, so that the migration halts at longer orbital periods. The bottom panel (blue histogram) shows the results from the simulations that began with the more distant, and wider spaced, initial conditions discussed in § 5.2.1. This means that more systems get trapped into the 2:1 resonance instead of the 3:2 resonances, and the later divergence leads to wider systems. These simulations assume the standard $a_{R}$.
@@
The other panels in Figure 14 show the output from our simulations. The second panel from the top, labelled as ‘Original’, represents the results from our default simulations (the ones shown as red points in Figure 10). These have been assigned weights according to their geometric probability of observation, as discussed above. The third panel from the top represents the evolution of the same set of initial conditions, but with $a_{R}$ scaled up by a factor of 1.2 relative to the default value given in equation (4). This represents the case of a stronger stellar magnetic fields. The resulting distribution is pretty similar, although the resonant peaks are less pronounced – i.e. a larger fraction of systems moved away from commensurabilities.
@@
Finally, the lower panel shows the results of the simulations discussed in § 5.2.1. These start with more widely separated initial conditions (and an $a_{R}$ corresponding to that of the second panel, i.e. all systems have the default $a_{R}$). We see that many more of these systems get captured in the 2:1 resonance and we get a distribution that skews to wider separations than seen in the observations.
@@
A more quantitative comparison of the non-resonant fractions is shown in Figure 15, which compares the cumulative distribution of period ratios between theory and observations, using the same colors as in Figure 14.
@@
Thus, the black histogram represents the observed distribution drawn from the current Kepler sample (the solid histogram in the upper panel Figure 14). The red histogram represents the original model sample from the second panel, and which reproduces the observed distribution quite accurately. The magenta histogram shows the distribution from the third panel. This one shows too many at lower period ratios and too few at higher period ratios. The blue histogram shows the sample from the lowest panel – those drawn from wide initial conditions. In this case the model clearly produces too few close pairs and even falls below the observed curve at larger periods, indicating a substantial pile-up of systems between period ratios of 2.2 and 3. Overall, we find that the default model does an excellent job of providing a distribution of period ratios broadly consistent with those observed. A Kolmogorov-Smirnov test indicates that the two distributions have a 70% chance of being drawn from the same underlying distribution. The observations can also be reproduced with combinations of the different samples. The cyan histogram shows a model in which 70% of the systems come from the simulations with larger $a_{R}$ and compact initial conditions, along with 30% of the systems that began with wider initial conditions.
@@
Another useful comparison set is the simulations of Izidoro et al. (2017) and Izidoro et al. (2021), the most comprehensive effort to date to compare the migration model with observations in a quantitative fashion. Izidoro et al. (2017) describe a model for the migration of low mass planets in both static and turbulent disks, which decay away on Myr timescales, leaving planets trapped in a chain of mean motion resonances. They furthermore run the simulations forward another 100 Myr after the dispersal of the gas disk, and track the evolution of systems that undergo dynamical instabilities which can break the resonant chains. The encouraging part of the results of Izidoro et al. (2017) is that the resulting instabilities do indeed produce many non-resonant systems. However, the fraction of resonant systems that remain is still far too high ($\sim 50\%$) and the dynamical instabilities lead to planetary collisions which result in systems too massive compared to the observed systems. The simulations of Izidoro et al. (2021) takes these models further by including a prescription for gas-assisted pebble accretion during the migration phase. Variation in the pebble flux and the formation time of planetesimals result in various levels of dynamical instability. The authors find that it is possible to match the neighbour period ratio distribution of the Kepler planets but, once again, only if one treats the fraction of unstable systems as a free parameter. For some parameter choices, most of the simulated systems do go unstable, but these also, once again, produce systems that are too massive.
@@
For comparison, the dotted histogram shows the outcomes from the in situ assembly simulations of Hansen & Murray (2013) for the case of a total solid inventory of $20M_{\oplus}$. Both distributions show a broad spread around a single peak, displaced by a factor of a few from the nominal stability limit (Pu & Wu (2015) and references therein). The magnetospheric rebound model actually produces systems that are even more compact than the in situ models, despite the fact that the latter are partially sculpted by considerations of dynamical stability, while the spacing of the former is regulated by the freeze-out process of the gaseous disk. As a consequence, it is misleading to infer that a closely packed planetary system is automatically the result of sculpting by late-stage dynamical instabilities.
The time at which a given planet freezes out of the migration will also determine the final orbital period of the planet. This will also be affected by the value of $a_{R}$ and hence the stellar magnetic field. To test whether our model values are reasonable, we simulated the same 107 Kepler systems with the default value of $a_{R}$, a version with $a_{R}$ scaled up by a factor 1.2 (discussed above already) and also one with a scaling factor of 0.8. Figure 17 shows the resulting distribution of final planetary locations for all planets summed over these three simulations. In this case we do not adjust the counts for the geometric probabilities because the observations to which we compare have already accounted for selection effects and purport to represent the true underlying distribution. The shaded regions are these corresponding observed trends in planet occurrence for Super-Earths, as measured by Petigura et al. (2018). The simulations provide an excellent match to the observed rise from short periods to a peak in the range of $\sim$ 6–8 days, followed by a drop-off towards larger periods. Immediately exterior to the peak, the simulations also match the slope measured by Petigura et al. (2018) but begin to drop below this at orbital periods $>$20 days. This is a reflection of our restriction to triple systems. Systems with additional planets are expected to further fill in the distribution at longer periods.
@@ 7 Conclusions
The results presented here demonstrate that a detailed treatment of the inner boundary of protoplanetary disks (Yu et al., 2023) has a strong influence on the final configurations of planetary systems that migrate through this disk. In particular, these models show the existence of an inner transition region of outwardly directed migration, which causes divergent migration of planets in resonant chains at late times. This results in a substantial fraction of initially resonant chains diverging to the point that their period ratios no longer lie close to first order commensurabilities (although some may still be technically resonant because of their high precession rates). The resulting distribution of planetary neighbour period ratios is a good match to the observed distribution of planets detected by Kepler, and also provides a good match to the observed relative occurrence frequency of observed Super-Earth planets. These results provide an explanation for the long-standing problem that the observed distribution of period ratios showed far fewer resonant systems than expected in the models.
@@
The authors thank the referees for prompt and comprehensive reports, which improved the exposition. This research has made use of the NASA Exoplanet Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. This research has made use of NASA’s Astrophysics Data System Bibliographic Services.
@@ Appendix A A: Dynamical Equations
The planetary dynamics will be described in terms of the orbital frequencies, $n_{1}$, $n_{2}$ and $n_{3}$, normalised to a fidicial value $n_{0}$, so that $x_{i}=n_{i}/n_{0}$. These correspond to semi-major axis $a_{1}<a_{2}<a_{3}$, so that $\alpha_{12}=a_{1}/a_{2}$ and $\alpha_{23}=a_{2}/a_{3}$. Eccentricities are denoted by $e_{i}$ and the arguments of periastron are designed as $\bar{\omega}_{i}$. We will incorporate effects of the first order resonances $(1+q)/q$ for $q=1$ and $q=2$, between nearest neighbours. The commensurate angles are $\psi_{1}=2\lambda_{2}-\lambda_{1}$, $\psi_{2}=3\lambda_{2}-2\lambda_{1}$, $\psi_{3}=2\lambda_{3}-\lambda_{2}$ and $\psi_{4}=3\lambda_{3}-2\lambda_{2}$. We normalise the timescale to $t_{0}=M_{*}/(3n_{0}m_{2})$, so that $\dot{x}_{i}=dx_{i}/d\tau=t_{0}dx_{i}/dt$. The resonant angles are constructed by combining these commensurate angles with the relevant longitudes of periastron.
@@
$\dot{x}_{1}$ $=$ $-x_{1}^{2}\alpha_{12}\left[A_{12}\sin\psi_{1}-B_{12}\cos\psi_{1}\right]-2x_{1}^{2}\alpha_{12}\left[C_{12}\sin\psi_{2}-D_{12}\cos\psi_{2}\right]+\frac{3}{2}\frac{x_{1}}{\tau_{a,1}}+3e_{1}^{2}\frac{x_{1}}{\tau_{e,1}}$ (7)
$\dot{x}_{2}$ $=$ $2x_{2}^{2}\frac{m_{1}}{m_{2}}\left[A_{12}\sin\psi_{1}-B_{12}\cos\psi_{1}\right]+3x_{2}^{2}\frac{m_{1}}{m_{2}}\left[C_{12}\sin\psi_{2}-D_{12}\cos\psi_{2}\right]+$ (8)
$-x_{2}^{2}\alpha_{23}\frac{m_{3}}{m_{2}}\left[A_{23}\sin\psi_{3}-B_{23}\cos\psi_{3}\right]-2x_{2}^{2}\alpha_{23}\frac{m_{3}}{m_{2}}\left[C_{23}\sin\psi_{4}-D_{23}\cos\psi_{4}\right]+\frac{3}{2}\frac{x_{1}}{\tau_{a,1}}+3e_{2}^{2}\frac{x_{2}}{\tau_{e,2}}$
$\dot{x}_{3}$ $=$ $2x_{3}^{2}\left[A_{23}\sin\psi_{3}-B_{23}\cos\psi_{3}\right]+3x_{3}^{2}\left[C_{23}\sin\psi_{4}-D_{23}\cos\psi_{4}\right]+\frac{3}{2}\frac{x_{3}}{\tau_{a,3}}+3e_{3}^{2}\frac{x_{3}}{\tau_{e,3}}$ (9)
@@
The quantities $\tau_{a,i}$ represent the eccentricity damping timescales given by equation (2.1), normalised to $t_{0}$. The quantities $\tau_{e,i}$ represent the corresponding eccentricity damping timescales, taken to be $\tau_{e,i}=0.03\left|\tau_{a,i}\right|$ (after fitting to the calculations of Yu et al. (2023)). Eccentricity damping for a given planet $i$ is switched off if $e_{i}<10^{-4}$. This eccentricity damping can often lead to low eccentricities which, in turn, leads to high precession rates. As a result, some of the resonant angles can continue to librate even when the period ratios deviate from the exact commensurability. Our equations represent the expansions in terms of the resonant arguments for the instantaneous osculating value of $\alpha$, and so we describe the functions $f_{j}$ as functions of $\alpha$ and not simply a number at exact commensurability.
@@
$\dot{h}_{1}$ $=$ $-\frac{h_{1}}{\tau_{e,1}}+\frac{\alpha_{12}}{3}f_{1}(\alpha_{12})x_{1}\cos\psi_{1}+\frac{\alpha_{12}}{3}f_{3}(\alpha_{12})x_{1}\cos\psi_{2}$ (20)
$\dot{k}_{1}$ $=$ $-\frac{k_{1}}{\tau_{e,1}}-\frac{\alpha_{12}}{3}f_{1}(\alpha_{12})x_{1}\sin\psi_{1}+\frac{\alpha_{12}}{3}f_{3}(\alpha_{12})x_{1}\sin\psi_{2}$ (21)
$\dot{h}_{2}$ $=$ $-\frac{h_{2}}{\tau_{e,2}}+\frac{1}{3}\frac{m_{1}}{m_{2}}f^{\prime\prime}_{2}(\alpha_{12})x_{2}\cos\psi_{1}+\frac{1}{3}\frac{m_{1}}{m_{2}}f_{4}(\alpha_{12})x_{2}\cos\psi_{2}$ (22)
$+\frac{\alpha_{23}}{3}\frac{m_{3}}{m_{2}}f_{1}(\alpha_{23})\cos\psi_{3}+\frac{\alpha_{23}}{3}\frac{m_{3}}{m_{2}}f_{3}(\alpha_{23})x_{2}\cos\psi_{4}$
$\dot{k}_{2}$ $=$ $-\frac{k_{2}}{\tau_{e,2}}-\frac{1}{3}\frac{m_{1}}{m_{2}}f^{\prime\prime}_{2}(\alpha_{12})x_{2}\sin\psi_{1}-\frac{1}{3}\frac{m_{1}}{m_{2}}f_{4}(\alpha_{12})x_{2}\sin\psi_{2}$ (23)
$-\frac{\alpha_{23}}{3}\frac{m_{3}}{m_{2}}f_{1}(\alpha_{23})x_{2}\sin\psi_{3}+\frac{\alpha_{23}}{3}\frac{m_{3}}{m_{2}}f_{3}(\alpha_{23})x_{2}\sin\psi_{4}$
$\dot{h}_{3}$ $=$ $-\frac{h_{3}}{\tau_{e,3}}+\frac{1}{3}f^{\prime\prime}_{2}(\alpha_{23})x_{3}\cos\psi_{3}+\frac{1}{3}f_{4}(\alpha_{23})x_{2}\cos\psi_{4}$ (24)
$\dot{k}_{3}$ $=$ $-\frac{k_{3}}{\tau_{e,3}}-\frac{1}{3}f^{\prime\prime}_{2}(\alpha_{23})x_{3}\sin\psi_{3}-\frac{1}{3}f_{4}(\alpha_{23})x_{3}\sin\psi_{4}$ (25)
Finally, the above equations depend on the commensurabilities $\psi_{1}=2\lambda_{2}-\lambda_{1}$, $\psi_{2}=3\lambda_{2}-2\lambda_{1}$, $\psi_{3}=2\lambda_{3}-\lambda_{2}$ and $\psi_{4}=3\lambda_{3}-2\lambda_{2}$. The evolution of these variables is given by
$\dot{\psi}_{1}$ $=$ $\frac{1}{\Delta}\left(2x_{2}-x_{1}\right)$ (26)
$\dot{\psi}_{2}$ $=$ $\frac{1}{\Delta}\left(3x_{2}-2x_{1}\right)$ (27)
$\dot{\psi}_{3}$ $=$ $\frac{1}{\Delta}\left(2x_{3}-x_{2}\right)$ (28)
$\dot{\psi}_{4}$ $=$ $\frac{1}{\Delta}\left(3x_{3}-2x_{2}\right)$ (29)
where $\Delta=3m_{2}/M_{*}$. The true resonant angles $\phi$ are determined by adding the relevant longitudes of periastron to the corresponding commensurabilities.